If i have a rectangle with an area of 66 sq cm, and a length of 13 cm less than four times the width, how do i find its dimensions?
Answer by checkley75(3666)

L*W=AREA
LENGTH=4W-13 & WIDTH=W. THEN
(4W-13)W=66
4W^2-13W-66=0
(4W+11)(W-6)=0
4W+11=0
4W=-11
W=-11/4 NOT AN ANSWER.
W-6=0
W=6 ANSWER FOR THE WIDTH.
L=4*6-12
L=24-13
L=11 ANSWER.
PROOF
6*11=66
66=66